# Kernel Density Steering: Inference-Time Scaling via Mode Seeking for Image Restoration

### v1

@@ 1 Introduction
This paper introduces Kernel Density Steering (KDS), an inference-time framework designed to improve the distortion-perception trade-off without requiring any model retraining. KDS leverages collective information from an $N$-particle ensemble by steering particles towards regions of high concentration within the ensemble. To identify these regions, KDS employs Kernel Density Estimation (KDE) [35, 36]. Specifically, at each time step t, we estimate a smoothed probability density $\hat{p}_{K}({\bm{z}}_{t}|{\bm{c}})$ over the particle ensemble via KDE:
$\hat{p}_{K}({\bm{z}}_{t}|{\bm{c}})=\frac{1}{Nh^{d}}\sum_{k=1}^{N}K\left(\frac{\|{\bm{z}}_{t}-{\bm{z}}_{t}^{(k)}\|^{2}}{h^{2}}\right),$ (1)
where $h$ is bandwidth, $d$ is dimensionality. KDS then steers particles towards the modes of this dynamically estimated density $\hat{p}_{K}$. We hypothesize these modes, areas of strong particle agreement, yield solutions with a better distortion-perception balance. The steering direction for each particle ${\bm{z}}_{t}^{(i)}$ is the gradient of this log-density: $\nabla{{\bm{z}}_{t}}{\mathsf{log}}\hat{p}_{K}({\bm{z}}_{t}^{(i)}|{\bm{c}})$, which points towards the steepest increase in $\hat{p}_{K}({\bm{z}}_{t}|{\bm{c}})$. This KDS guidance is added to the standard diffusion sampling update for each particle $i$:
$d{\bm{z}}_{t}^{(i)}=\left[\underbrace{f(t){\bm{z}}_{t}^{(i)}-\frac{1}{2}g(t)^{2}\nabla_{{\bm{z}}_{t}}{\mathsf{log}}p_{t}({\bm{z}}_{t}^{(i)}|{\bm{c}})}_{\text{Standard Diffusion Update}}\right]dt+\underbrace{\delta_{t}g(t)^{2}\nabla_{{\bm{z}}_{t}}{\mathsf{log}}\hat{p}_{K}({\bm{z}}_{t}^{(i)}|{\bm{c}})}_{\text{Kernel Density Steering Term}}dt,$ (2)
where $\delta_{t}$ is the KDS guidance strength. Thus, particles are influenced by the original model’s score and concurrently guided by the KDS term, which acts as a form of collaborative filtering across the multiple particles, pushing them towards areas of high ensemble consensus. This adaptive, internal guidance aims for restorations that are both quantitatively accurate and perceptually sharp.
@@
Figure 1: Conceptual illustration of Kernel Density Steering (KDS) for diffusion sampling. Left: Standard diffusion sampling with $N$ independent particles (blue dots) can result in high variance. Right: KDS utilizes the ensemble of $N$ particles to estimate local density. It then guides these particles towards shared high-density modes (peaks in the density curve) during the diffusion process.
@@ 2.1 Latent Diffusion Models
To sample from the learned distribution, Denoising Diffusion Implicit Models (DDIM) [38] offer a fast, non-Markovian sampling procedure. The sampling step from ${\bm{z}}_{t}$ to ${\bm{z}}_{t-1}$ in DDIM, which predicts the latent state at $t-1$ based on the estimated clean latent at time $t$, is given by:
@@ 2.2 LDM for Image Restoration
Two primary methods exist for leveraging diffusion models in IR. Conditional LDMs [4, 5, 6, 9, 10, 11, 15, 16, 18, 19, 20, 39], learn the posterior $p({\bm{z}}|{\bm{c}})$ directly, where the conditioning ${\bm{c}}$ is derived from ${\bm{y}}$. These models can be highly effective but may face challenges in accurately modeling the true posterior for complex degradations, potentially impacting restoration fidelity. Alternatively, unconditional LDMs can be employed with inference-time guidance using Bayes’ rule  [40, 41, 42, 43, 44]. These methods typically require computing gradients of the log-likelihood ${\mathsf{log}}p({\bm{y}}|{\bm{x}}_{t})$ (often approximated via the estimate $\hat{{\bm{x}}}_{0|t}$). This necessitates knowledge of the degradation process (e.g., the operator $A$ in ${\bm{y}}\approx A{\bm{x}}$), limiting their applicability in scenarios where the degradation is unknown or cannot be accurately modeled which is the case in many real-world image SR or deblurring.
@@ 3 Kernel Density Steering for Diffusion Samplers
KDS achieves this ensemble-driven mode-seeking by incorporating an additional steering term into the sampling update for each particle. This steering is inspired by the Mean Shift algorithm (Sec. 2.4). Specifically, for each particle, KDS computes a mean shift vector $\mathbf{m}({\bm{z}}_{t}^{(i)})$ based on the current ensemble ${\bm{Z}}_{t}$. Then KDS steering is then applied by taking a step in the direction of this mean shift vector, which directs to a local mode of the ensemble’s Kernel Density Estimate (KDE).
@@ 4.1 KDS Sampling in a 2D Example Case
+ KDS ($N=10$) 22.79 0.587 0.313 5.026 0.191 0.488 0.679 25.44
@@ 4.2 Real-world Image Restoration
Figure 3: Qualitative comparison of 4×4$\times$4 × Real-world Super-Resolution with KDS-enhanced DDIM sampling (Number of Particles $N=10$). ’w/o KDS’ shows results from baseline DDIM, while ’w/ KDS’ shows results with KDS. KDS consistently produces images with improved sharpness, finer details, and reduced artifacts across LDM-SR, DiffBIR, and SeeSR backbones.
@@
+ KDS ($N=10$) 24.50 0.719 0.272 0.206 0.640 27.42 0.765 0.287 0.212 0.651
@@ 4.2.1 Real-world Image Super-Resolution
We evaluated the performance of KDS for 4×4$\times$4 × real-world super-resolution (SR). This evaluation utilized the LDM-SR [24], DiffBIR [4], and SeeSR [33] backbones across several datasets: DIV2K [55], RealSR [56], DrealSR [57]. Performance was evaluated using a comprehensive suite of metrics, including PSNR, SSIM [58], LPIPS [59], FID [60], NIMA [61], MANIQA [62], and CLIPIQA [63].
@@
Compare with Best-of-N approaches: While inference-time scaling allows for the generation of multiple candidate solutions, especially beneficial for low-quality inputs, selecting the optimal one from an N-particle ensemble is non-trivial. As Table 4 shows, using a non-reference metric like LIQE [64] to pick the best particle results in significantly lower performance compared to our Kernel Density Smoothing (KDS) method. This demonstrates that, despite comparable computational costs, traditional post-sampling selection methods do not achieve the same level of performance as KDS.
@@
+ KDS ($N\!=\!10$) 21.35 0.748 0.131 11.18
Table 4: Comparison to Best-of-$N$ (BoN) using LIQE [64] on RealSR with DiffBIR.
Method $N$ PSNR LPIPS ($\downarrow$) Time Memory
@@ 4.3 Analysis of KDS Robustness
Figure 6: Influence of Particle Number ($N$) and Bandwidth ($h$) on Real-world SR. Performance metrics (PSNR, SSIM, LPIPS, FID) are plotted against the number of particles ($N$) for different bandwidth ($h$) settings on the DIV2K dataset, using LDM-SR backbone with KDS.
@@ 4.4 Impact of KDS Hyperparameters: Number of Particles and Bandwidth
Kernel Bandwidth $\bm{h}$: The kernel bandwidth $h$ plays a crucial role in balancing the perception-distortion trade-off. Figure 6 (different colored lines representing different $h$ settings) demonstrates that while larger ensembles generally improve performance, the choice of $h$ is critical. Excessively large bandwidths (e.g., $h=0.6$ in some cases) can lead to over-smoothing, potentially enhancing distortion metrics like PSNR & SSIM at the cost of perceptual quality (higher LPIPS & FID). Conversely, a very small $h$ might not capture sufficient local consensus. Optimal $h$ selection is therefore essential for achieving the desired balance between quantitative accuracy and perceptual sharpness. Our experiments suggest that a moderate $h$ often provides a good compromise.
@@ 5 Conclusion
Limitations & Future Work. Kernel Density Steering (KDS), like other ensemble methods, introduces a computational overhead that scales linearly with the number of particles. While our results indicate that moderate ensemble sizes (e.g., $N=10-15$) offer a good balance between performance and cost, further research could explore strategies to mitigate this overhead, such as adaptive ensemble sizes or more efficient kernel density estimation techniques. Additionally, the performance of KDS is influenced by the choice of kernel bandwidth ($h$), which requires careful tuning to avoid over-smoothing or insufficient consensus guidance. Future work could investigate adaptive bandwidth selection methods to further optimize performance across diverse image restoration tasks.
@@ Acknowledgements
The authors would like to thank our colleagues Keren Ye and Mo Zhou for reviewing the manuscript and providing valuable feedback.
@@ A.1.3 Patch-wise KDS Mechanism: Configuration and Discussion
Patch Size Configuration: To effectively implement our patch-wise strategy, the patch size is chosen to balance accurate mode-seeking with computational efficiency. Accurate mode-seeking across the entire latent space would necessitate throunds of samples, which is impractical for real-world applications. Therefore, to achieve robust estimation with minimal samples, we set the patch size to $1\times 1$. This processes each spatial location $(h,w)$ in the latent map as an individual $8$-dimensional vector (i.e., batches of $1\times 1\times 8$ vectors). This strategic choice significantly reduces the dimensionality for each kernel density estimation (KDE), enabling accurate estimation with limited samples while leveraging the rich, learned features of the $8$ channels. Consequently, this facilitates more effective mode-seeking with a limited particle count. Ablation studies (Table 5) demonstrate that with a restricted particle count ($N=10$), larger patch sizes lead to inaccurate mode-seeking and sub-optimal performance.
@@ A.1.4 Integration with Standard Diffusion Samplers
As discussed in the main paper, KDS is a flexible, plug and play approach that can be applied to all existing samplers. In this section, we will detailed introduce how to apply KDS to DDIM, and DPM-solver and how to interatve with Classifier-Free Guidance.
@@
4:   $\bm{\epsilon}_{\theta,t}\leftarrow\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}})$triangleright Predict noise
5:   $\hat{{\bm{z}}}_{0|t}\leftarrow({\bm{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta,t})/\sqrt{\bar{\alpha}_{t}}$triangleright Predict ${\bm{z}}_{0}$
6:   $\bm{\epsilon}^{\prime}_{t}\leftarrow({\bm{z}}_{t}-\sqrt{\bar{\alpha}_{t}}\hat{{\bm{z}}}_{0|t})/\sqrt{1-\bar{\alpha}_{t}}$triangleright Update direction based on $\hat{{\bm{z}}}_{0|t}$
7:   ${\bm{z}}_{t-1}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\hat{{\bm{z}}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}\bm{\epsilon}^{\prime}_{t}$triangleright DDIM step
@@
1:Model $\bm{\epsilon}_{\theta}$, Schedule $\bar{\alpha}_{t}$, condition: ${\bm{c}}$, Number of particles: $N$, bandwidth: $h$, steering strength: $\delta_{t}$, PatchSize
2:${\bm{Z}}_{T}\sim\mathcal{N}(0,\mathbf{I})$ (ensemble of $N$ samples) $\triangleright$ Init noise ensemble
@@
4:    $\mathbf{E}_{\theta,t}\leftarrow\bm{\epsilon}_{\theta}({\bm{Z}}_{t},t,{\bm{c}})$triangleright Predict ensemble noise
5:    $\hat{{\bm{Z}}}_{0|t}^{\text{pred}}\leftarrow({\bm{Z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{E}_{\theta,t})/\sqrt{\bar{\alpha}_{t}}$triangleright Predict $\mathbf{x}_{0}$ ensemble
6:    $\textbf{Patches}\leftarrow\text{Patchify}(\hat{{\bm{Z}}}_{0|t}^{\text{pred}})$triangleright Extract all non-overlapped patches: $\textbf{Patches}[k,loc]$.
@@
8:       ${\bm{P}}_{j}\leftarrow\textbf{Patches}[:,j]$triangleright Ensemble of $N$ original patches at location $j$.
@@
10:          ${\bm{p}}_{j}^{(i)}\leftarrow{\bm{P}}_{j}[i]$triangleright Patch from particle $i$ at location $j$.
11:          $\mathbf{m}({\bm{p}}_{j}^{(i)})\leftarrow\frac{\sum_{k=1}^{N}G\left(\frac{\|{\bm{p}}_{j}^{(i)}-{\bm{P}}_{j}^{(k)}\|^{2}}{h^{2}}\right){\bm{P}}_{j}^{(k)}}{\sum_{k=1}^{N}G\left(\frac{\|{\bm{p}}_{j}^{(i)}-{\bm{P}}_{j}^{(k)}\|^{2}}{h^{2}}\right)}-{\bm{p}}_{j}^{(i)}$triangleright Mean shift vector (Eq. 7).
12:          $\hat{{\bm{p}}}_{j}^{(i),\text{KDS}}\leftarrow{\bm{p}}_{j}^{(i)}+\delta_{t}\mathbf{m}({\bm{p}}_{j}^{(i)})$triangleright Apply steering (Eq. 8)
13:          $\textbf{Patches}[i,j]\leftarrow\hat{{\bm{p}}}_{j}^{(i),\text{KDS}}$triangleright Update the patch set with the guided patch.
@@
16:    $\hat{{\bm{Z}}}_{0|t}^{\text{KDS}}\leftarrow\text{Unpatchify}(\textbf{Patches})$triangleright Reconstruct guided latent prediction.
17:    $\mathbf{E}^{\prime}_{t}\leftarrow({\bm{Z}}_{t}-\sqrt{\bar{\alpha}_{t}}\hat{{\bm{Z}}}_{0|t}^{\text{KDS}})/\sqrt{1-\bar{\alpha}_{t}}$triangleright Update direction
18:    ${\bm{Z}}_{t-1}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\hat{{\bm{Z}}}_{0|t}^{\text{KDS}}+\sqrt{1-\bar{\alpha}_{t-1}}\mathbf{E}^{\prime}_{t}$triangleright DDIM step
@@
20:return $\hat{{\bm{Z}}}_{0}^{\text{KDS}}$triangleright Return KDS-guided result
@@
1:Model $\bm{\epsilon}_{\theta}$, Schedule $\bar{\alpha}_{t}$, condition: ${\bm{c}}$, Number of particles: $N$, bandwidth: $h$, steering strength: $\delta_{t}$, CFG strength: $w$, PatchSize
2:${\bm{Z}}_{T}\sim\mathcal{N}(0,\mathbf{I})$ (ensemble of $N$ samples) $\triangleright$ Init noise ensemble

### v2

@@ 1 Introduction
This paper introduces Kernel Density Steering (KDS), an inference-time framework designed to improve the distortion-perception trade-off without requiring any model retraining. KDS leverages collective information from an $N$-particle ensemble by steering particles towards regions of high concentration within the predicted solution ensemble.
This mechanism is implemented as a two-step process. First, at each step $t$, the denoiser predicts the clean data $\hat{\mathbf{z}}_{0}^{(i)}$ for each particle $i$. Second, it performs a non-parametric mode-seeking operation to steer each prediction $\hat{\mathbf{z}}_{0}^{(i)}$ toward a local region of high density. This is achieved by implicitly estimating the ensemble’s density using Kernel Density Estimation (KDE) [35, 36] and then applying a single mean-shift update—which is equivalent to a gradient ascent step on this estimated density—to find a consensus point $\tilde{\mathbf{z}}_{0}^{(i)}$:
$\tilde{\mathbf{z}}_{0}^{(i)}=\frac{\sum_{k=1}^{N}K\left(\frac{\|\hat{\mathbf{z}}_{0}^{(i)}-\hat{\mathbf{z}}_{0}^{(k)}\|^{2}}{h^{2}}\right)\hat{\mathbf{z}}_{0}^{(k)}}{\sum_{k=1}^{N}K\left(\frac{\|\hat{\mathbf{z}}_{0}^{(i)}-\hat{\mathbf{z}}_{0}^{(k)}\|^{2}}{h^{2}}\right)}.$ (1)
This refined estimate $\tilde{\mathbf{z}}_{0}^{(i)}$ defines a new consensus score, $\tilde{\mathbf{s}}_{t}^{(i)}$. The original and consensus scores are defined as:
$\mathbf{s}_{t}^{(i)}=\frac{\mathbf{z}_{t}^{(i)}-\sqrt{\bar{\alpha}_{t}}\hat{\mathbf{z}}_{0}^{(i)}}{\sigma_{t}^{2}}\quad\text{and}\quad\tilde{\mathbf{s}}_{t}^{(i)}=\frac{\mathbf{z}_{t}^{(i)}-\sqrt{\bar{\alpha}_{t}}\tilde{\mathbf{z}}_{0}^{(i)}}{\sigma_{t}^{2}}.$ (2)
This guidance is incorporated into the probability flow ODE via a weighted average of the two scores:
$d\mathbf{z}_{t}^{(i)}=\left[f(t)\mathbf{x}_{t}^{(i)}-\frac{1}{2}g(t)^{2}\left((1-\delta_{t})\mathbf{s}_{t}^{(i)}+\delta_{t}\tilde{\mathbf{s}}_{t}^{(i)}\right)\right]dt,$ (3)
where $\delta_{t}$ is the KDS guidance strength. Thus, particles are guided by the original score and the KDS term. This term functions as a form of collaborative filtering, using non-parametric mode-seeking to guide the ensemble toward areas of high consensus. This adaptive, internal guidance aims for restorations that are both quantitatively accurate and perceptually sharp.
@@
Figure 1: Conceptual illustration of Kernel Density Steering (KDS) for diffusion sampling. Left: Standard diffusion sampling with NN independent particles (blue dots) can result in high variance. Right: KDS utilizes the ensemble of $N$ particles to estimate local density. It then guides these particles towards shared high-density modes (peaks in the density curve) during the diffusion process.
@@ 2.1 Latent Diffusion Models
To sample from the learned distribution, Denoising Diffusion Implicit Models (DDIM) [38] offer a fast, non-Markovian sampling procedure. The sampling step from ${\bm{z}}_{t}$ to ${\bm{z}}_{t-1}$ in DDIM, which predicts the latent state at t1t-1 based on the estimated clean latent at time $t$, is given by:
@@ 2.2 LDM for Image restoration
Two primary methods exist for leveraging diffusion models in IR. Conditional LDMs [4, 5, 6, 9, 10, 11, 15, 16, 18, 19, 20, 39], learn the posterior $p({\bm{z}}|{\bm{c}})$ directly, where the conditioning $\bm{c}$} is derived from $\bm{y}$}. These models can be highly effective but may face challenges in accurately modeling the true posterior for complex degradations, potentially impacting restoration fidelity. Alternatively, unconditional LDMs can be employed with inference-time guidance using Bayes’ rule  [40, 41, 42, 43, 44, 45]. These methods typically require computing gradients of the log-likelihood ${\mathsf{log}}p({\bm{y}}|{\bm{x}}_{t})$ (often approximated via the estimate $\hat{{\bm{x}}}_{0|t}$). This necessitates knowledge of the degradation process (e.g., the operator $A$ in ${\bm{y}}\approx A{\bm{x}}$), limiting their applicability in scenarios where the degradation is unknown or cannot be accurately modeled which is the case in many real-world image SR or deblurring.
@@ 3 Kernel Density Steering for Diffusion Samplers
KDS achieves this ensemble-driven mode-seeking by incorporating an additional steering term into the sampling update for each particle. This steering is inspired by the Mean Shift algorithm (Sec. 2.4). Specifically, for each particle, KDS computes a mean shift vector $\mathbf{m}({\bm{z}}_{t}^{(i)})$ based on the current ensemble ${\bm{Z}}_{t}$. Then KDS steering is applied by taking a step in the direction of this mean shift vector, which directs to a local mode of the ensemble’s Kernel Density Estimate (KDE).
@@ 4.1 KDS Sampling in a 2D Example Case
[…]
Statistically significant differences ($P<0.05$) compared with DDIM are marked with a $\star$.
@@ 4.2 Real-world Image Restoration
Figure 3: Qualitative comparison of $4\times$ Real-world Super-Resolution with KDS-enhanced DDIM sampling (Number of Particles $N=10$). ’w/o KDS’ shows results from baseline DDIM, while ’w/ KDS’ shows results with KDS. KDS consistently produces images with improved sharpness, finer details, and reduced artifacts across LDM-SR, DiffBIR, and SeeSR backbones.
@@
[…]
Statistically significant differences ($P<0.05$) compared with DDIM are marked with a $\star$.
@@ 4.2.1 Real-world Image Super-Resolution
We evaluated the performance of KDS for $4\times$ real-world super-resolution (SR). This evaluation utilized the LDM-SR [24], DiffBIR [4], and SeeSR [33] backbones across several datasets: DIV2K [56], RealSR [57], DrealSR [58]. Performance was evaluated using a comprehensive suite of metrics, including PSNR, SSIM [59], LPIPS [60], FID [61], NIMA [62], MANIQA [63], and CLIPIQA [64].
@@
Compare with Best-of-N approaches: While inference-time scaling allows for the generation of multiple candidate solutions, especially beneficial for low-quality inputs, selecting the optimal one from an N-particle ensemble is non-trivial. As Table 4 shows, using a non-reference metric like LIQE [65] to pick the best particle results in significantly lower performance compared to our kernel density steering (KDS) method. This demonstrates that, despite comparable computational costs, traditional post-sampling selection methods do not achieve the same level of performance as KDS.
@@
+ KDS (N=10N\!=\!10) 21.35 0.748 0.131 11.18
Table 4: Comparison to Best-of-NN (BoN) using LIQE [65] on RealSR with DiffBIR.
Method NN PSNR LPIPS ($\downarrow$) Time Memory
@@ 4.3 Analysis of KDS Robustness
Figure 6: Influence of Particle Number (NN) and Bandwidth ($h$) on Real-world SR. Performance metrics (PSNR, SSIM, LPIPS, FID) are plotted against the number of particles (NN) for different bandwidth ($h$) settings on the DIV2K dataset, using LDM-SR backbone with KDS.
@@ 4.4 Impact of KDS Hyperparameters: Number of Particles and Bandwidth
Kernel Bandwidth $\bm{h}$: The kernel bandwidth $h$ plays a crucial role in balancing the perception-distortion trade-off. Figure 6 (different colored lines representing different $h$ settings) demonstrates that while larger ensembles generally improve performance, the choice of $h$ is critical. Excessively large bandwidths (e.g., $h=0.6$ in some cases) can lead to over-smoothing, potentially enhancing distortion metrics like PSNR & SSIM at the cost of perceptual quality (higher LPIPS & FID). Conversely, a very small $h$ might not capture sufficient local consensus. Optimal $h$ selection is therefore essential for achieving the desired balance between quantitative accuracy and perceptual sharpness. Our experiments suggest that a moderate $h$ often provides a good compromise. Furthermore, in Appendix A.2, we show that standard automatic bandwidth selection strategies (e.g., median heuristic) achieve performance nearly identical to our manually tuned $h$. This demonstrates that KDS can be deployed effectively without tedious manual tuning, enhancing its practical applicability.
Further parameter analyses in Appendix A.2 confirm KDS is robust to hyperparameter choices, including different kernel functions (e.g., Gaussian, Laplacian). We also explore the trade-off between particle count ($N$) and patch size. As shown in the appendix, increasing $N$ enables the use of larger patch sizes, which enhances performance by better mitigating the curse of dimensionality.
5 Discussion
We further analyze KDS by comparing it to other inference-time techniques. We compare it against other mode-seeking sampling strategies (HDS [66], DGS [67]) and separately against other methods, like InDI [29], which are also designed to improve the perception-distortion trade-off.
Comparison with HDS [66] and DGS [67]. To position KDS relative to other inference-time mode-seeking techniques, we compare it against High Density Sampling (HDS) [66] and Density-Guided Sampling (DGS) [67]. We implemented both methods with the DiffBIR [4] backbone and swept their key hyperparameters (time $t$ for HDS, quantile $q$ for DGS) to find their best-performing configurations on the RealSR dataset, with a full parameter sweep available in Appendix A.2.
As shown in Table 5, HDS achieves high PSNR at a severe cost to perceptual quality (LPIPS, CLIPIQA), while DGS offers moderate improvement. In contrast, KDS ($N=10$) provides the best balance, uniquely achieving both a strong PSNR and the best perceptual scores (LPIPS, DISTS, CLIPIQA), confirming its superior perception-distortion trade-off.
The core conceptual difference lies in their assumptions. Both HDS [66] and DGS [67] are mode-seeking but rely on different assumptions compared with KDS. HDS assumes a unimodal Gaussian distribution, which is often violated by highly multi-modal IR posteriors. DGS, while non-parametric in a different sense, guides a single sample by progressively sharpening the marginal density $p_{t}(x_{t})$ at each step, forcing it toward a single high-density mode. In sharp contrast, KDS makes no explicit assumption about the posterior’s shape. It uses a non-parametric KDE on an ensemble of outputs ($\hat{{\bm{z}}}_{0|t}$), making it better suited to finding a consensus within the multi-modal posteriors common in IR tasks.
Comparison with InDI [29]. We also distinguish KDS from InDI [29], which also adjusts the distortion-perception trade-off at inference time. Their core approaches differ: InDI is not a posterior mode-seeking algorithm. It performs progressive refinement to the final estimation ${\bm{z}}_{0}$, which does not aim to sample from the posterior mode. KDS, in contrast, directly seeks a consensus mode from the final clean output ensemble ($\hat{{\bm{z}}}_{0}$). Moreover, KDS is a training-free, plug-and-play module, whereas InDI requires an integrated training and sampling pipeline.
Table 5: Comparison with other Mode-Seeking Methods on RealSR dataset with DiffBIR backbone.
Method PSNR $\pm\sigma$ SSIM $\pm\sigma$ LPIPS ($\downarrow$) $\pm\sigma$ DISTS ($\downarrow$) $\pm\sigma$ CLIPIQA $\pm\sigma$
[…]
Statistically significant differences ($P<0.05$) compared with DDIM are marked with a $\star$.
6 Limitations & Future Work
Limitations. The primary limitation of KDS is its computational overhead, which scales linearly with particle count $N$. Although N=1015N=10-15 provides a good performance-cost trade-off, this overhead could be mitigated by future work on adaptive ensemble sizes or more efficient KDE techniques.
Future Work. As a training-free, plug-and-play framework, KDS is highly generalizable. Promising directions include applying it to other inverse problems (e.g., deblurring, denoising, medical imaging) and video restoration, where its patch-wise mechanism could be extended to 3D spatio-temporal patches to enhance temporal consistency.
@@ Acknowledgements
The authors would like to thank our colleagues Keren Ye, Viraj Shah, Ashwini Pokle and Mo Zhou for reviewing the manuscript and providing valuable feedback.
@@ A.1.3 Patch-wise KDS Mechanism: Configuration and Discussion
Patch Size Configuration: To effectively implement our patch-wise strategy, the patch size is chosen to balance accurate mode-seeking with computational efficiency. Accurate mode-seeking across the entire latent space would necessitate thousands of samples, which is impractical for real-world applications. Therefore, to achieve robust estimation with minimal samples, we set the patch size to $1\times 1$. This processes each spatial location $(h,w)$ in the latent map as an individual $8$-dimensional vector (i.e., batches of $1\times 1\times 8$ vectors). This strategic choice significantly reduces the dimensionality for each kernel density estimation (KDE), enabling accurate estimation with limited samples while leveraging the rich, learned features of the $8$ channels. Consequently, this facilitates more effective mode-seeking with a limited particle count. Ablation studies (Table 6) demonstrate that with a restricted particle count ($N=10$), larger patch sizes lead to inaccurate mode-seeking and sub-optimal performance.
@@ A.1.4 Integration with Standard Diffusion Samplers
As discussed in the main paper, KDS is a flexible, plug and play approach that can be applied to all existing samplers. In this section, we will detailed introduce how to apply KDS to DDIM, and DPM-solver and how to interactive with Classifier-Free Guidance.
@@
[…]
@@
1:Model $\bm{\epsilon}_{\theta}$, Schedule $\bar{\alpha}_{t}$, condition: $\bm{c}$}, Number of particles: NN, bandwidth: $h$, steering strength: $\delta_{t}$, PatchSize
[…]
@@
4:  $\mathbf{E}_{\theta,t}\leftarrow\bm{\epsilon}_{\theta}({\bm{Z}}_{t},t,{\bm{c}})$ $\triangleright$ Predict ensemble noise
5:  $\hat{{\bm{Z}}}_{0|t}^{\text{pred}}\leftarrow({\bm{Z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{E}_{\theta,t})/\sqrt{\bar{\alpha}_{t}}$ $\triangleright$ Predict $\mathbf{x}_{0}$ ensemble
6:  $\textbf{Patches}\leftarrow\text{Patchify}(\hat{{\bm{Z}}}_{0|t}^{\text{pred}})$ $\triangleright$ Extract all non-overlapped patches: $\textbf{Patches}[k,loc]$.
@@
8:   ${\bm{P}}_{j}\leftarrow\textbf{Patches}[:,j]$ $\triangleright$ Ensemble of NN original patches at location $j$.
@@
10:    ${\bm{p}}_{j}^{(i)}\leftarrow{\bm{P}}_{j}[i]$ $\triangleright$ Patch from particle $i$ at location $j$.
11:    $\mathbf{m}({\bm{p}}_{j}^{(i)})\leftarrow\frac{\sum_{k=1}^{N}G\left(\frac{\|{\bm{p}}_{j}^{(i)}-{\bm{P}}_{j}^{(k)}\|^{2}}{h^{2}}\right){\bm{P}}_{j}^{(k)}}{\sum_{k=1}^{N}G\left(\frac{\|{\bm{p}}_{j}^{(i)}-{\bm{P}}_{j}^{(k)}\|^{2}}{h^{2}}\right)}-{\bm{p}}_{j}^{(i)}$ $\triangleright$ Mean shift vector (Eq. 8).
12:    $\hat{{\bm{p}}}_{j}^{(i),\text{KDS}}\leftarrow{\bm{p}}_{j}^{(i)}+\delta_{t}\mathbf{m}({\bm{p}}_{j}^{(i)})$ $\triangleright$ Apply steering (Eq. 9)
13:    $\textbf{Patches}[i,j]\leftarrow\hat{{\bm{p}}}_{j}^{(i),\text{KDS}}$ $\triangleright$ Update the patch set with the guided patch.
@@
16:  $\hat{{\bm{Z}}}_{0|t}^{\text{KDS}}\leftarrow\text{Unpatchify}(\textbf{Patches})$ $\triangleright$ Reconstruct guided latent prediction.
17:  $\mathbf{E}^{\prime}_{t}\leftarrow({\bm{Z}}_{t}-\sqrt{\bar{\alpha}_{t}}\hat{{\bm{Z}}}_{0|t}^{\text{KDS}})/\sqrt{1-\bar{\alpha}_{t}}$ $\triangleright$ Update direction
18:  ${\bm{Z}}_{t-1}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\hat{{\bm{Z}}}_{0|t}^{\text{KDS}}+\sqrt{1-\bar{\alpha}_{t-1}}\mathbf{E}^{\prime}_{t}$ $\triangleright$ DDIM step
@@
20:return $\hat{{\bm{Z}}}_{0}^{\text{KDS}}$ $\triangleright$ Return KDS-guided result
@@
1:Model $\bm{\epsilon}_{\theta}$, Schedule $\bar{\alpha}_{t}$, condition: $\bm{c}$}, Number of particles: NN, bandwidth: $h$, steering strength: $\delta_{t}$, CFG strength: $w$, PatchSize
[…]